OpenStudy (anonymous):

Solve the system:] 3x+4y-z=0 x-2y+z=-8 4x+3y-2z=-4

OpenStudy (allank):

Make z the subject of the first equation z=3x+4y [a] Substitute it in the second x-2y+(3x+4y)=-8 x-2y+3x+4y=-8 4x+2y=-8 4x=-8-2y x=-2-1/2y Substitute x in [a] z=3(-2-1/2y)+4y z=-6-3/2y+4y z=2.5y-6 So now we have x and z written in terms of y. Substitute z and x in the third 4x+3y-2z=-4 4(-2-0.5y)+3y-2(2.5y-6)=-4 -8-2y+3y-5y+12=-4 -4y+4=-4 -4y=-8 y=2 Solve for x and z: x=-2-1/2y x=-2-1=-3 z=2.5y-6 z=2.5(2)-6=-1 Whoa, that was long. I do hope I didn't make a miscalculation. Do check the answers by plugging them into the original equations. Hope it helps.

OpenStudy (anonymous):

??? Thanks so much... I will

OpenStudy (allank):

you're welcome. i hope the ??? mean u understand :)

OpenStudy (anonymous):

yes dont worry the questions mark were there b4 you answer... thanks so much

OpenStudy (allank):

no prob. best of luck =)

OpenStudy (anonymous):

I do have a nother question can I ask you?

OpenStudy (allank):

sure

OpenStudy (allank):

i hope it's not on systems....can't do that calc again :)

OpenStudy (anonymous):

no lol

OpenStudy (allank):

cool.

OpenStudy (anonymous):

The sum of three numbers is 108 . The first number is 8 less than the second . The third number is 2 times the first. What are the numbers?

OpenStudy (allank):

I'll guide you through this, so you'll do most of the thinking. So, lets let the first number be, say, x. okay?

OpenStudy (anonymous):

ok

OpenStudy (allank):

nice. first: x "The first number is 8 less than the second" so the second can be written as...?

OpenStudy (anonymous):

x-8

OpenStudy (anonymous):

no Y-8

OpenStudy (allank):

not quite, but you have the idea. . the first number is 8 less than the second. meaning the second is greater than the first by 8

OpenStudy (allank):

thus the second is...?

OpenStudy (allank):

The sum of three numbers is 108 . The first number is 8 less than the second . The third number is 2 times the first. What are the numbers? I'm sure you figured out this one eventually. Let the first number be x. The second will be (x+8) "The first number is 8 less than the second" The third will be 2x. "The third number is 2 times the first" Now, the sum of the three is 108. So we have a nice equation. x+(x+8)+2x=108 Solve for x, then you can calculate the value of the second and third numbers. Best wishes.